OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

can someone help me please and thank you

OpenStudy (jdoe0001):

what's the slope of y = 2x + 2 ?

OpenStudy (anonymous):

idk the 2x

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

see the slope of that equation now?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

so it would be 2

OpenStudy (jdoe0001):

so the slope of this equation is 2 so any perpendicular line to it, will have a slope of NEGATIVE RECIPROCAL of that slope so if this equation line has a slope of 2, a perpendicular line equation slope will be RECIPROCAL of 2 1/2 NEGATIVE of that -1/2

OpenStudy (anonymous):

when I tried doing this problem by myself i got y = 2x + 6

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3 these are the answer i had to choose from i thought it would be the 1 one

OpenStudy (jdoe0001):

so now we know 2 things about that other line which is perpendicular to this one it has a slope of -1/2 and passes through the point ( 6 , 3) so let us use the "point-slope intercept" form, that is \(\bf y-y_1=m(x-x_1)\) \(\large \begin{array}{rrlll}\\ (x_1,y_1) = (6,3)\\\quad \\ \hline\\ \quad \\ &3&-\cfrac{1}{2}&& 6\\ y-&y_1=&m&(x-&x_1) \end{array}\)

OpenStudy (anonymous):

so would it be the 3rd one y= 1/2x +6

OpenStudy (jdoe0001):

well, we dunno, if you replace those values in the point-slope form, that is \(\large \begin{array}{rllll}\\ m = slope\\ (x_1,y_1) = (6,3)\\\quad \\ \hline\\ \quad \\ &3&-\cfrac{1}{2}&& 6\\ y-&y_1=&m&(x-&x_1) \end{array}\) how would the point-slope form look like?

OpenStudy (anonymous):

um idk y-3-1/2x (x-6) is that right or no ? lol

OpenStudy (jdoe0001):

\(\bf m = \color{green}{slope}\\ (\color{red}{x_1},\color{blue}{y_1}) = (6,3)\\\quad \\ \quad \\ y-y_1=m(x-x_1) \implies y-\color{blue}{3}=\color{green}{-\cfrac{1}{2}}(x-\color{red}{6})\) then just solve for "y" to get your equation :)

OpenStudy (anonymous):

oh okay well I did it on a piece of paper and I got y = -1/2x + ^ is that right

OpenStudy (anonymous):

y=-1/2x+6*

OpenStudy (jdoe0001):

yeap, that's the line "that's perpendicular to y = 2x + 2 and passes through the point (6, 3)"

OpenStudy (anonymous):

thank you so much

OpenStudy (jdoe0001):

yw